# Enhancing Score-Based Sampling Methods with Ensembles

### v1

@@ 3.2 Moderate-dimensional Problems
For low-dimensional problems ($D<10$), the linear operator $\boldsymbol{b}_{t}$ and the scaling matrix $\boldsymbol{g}_{t}$ in Equation (1) can be replaced with scalar terms such that the forward diffusion process evolves with equal variance in each dimension. As $D$ increases, isotropic random walks can cause ensemble members to move to areas of low probability, similarly to the standard Metropolis-Hastings algorithm. One mitigation strategy is to convert the SDE in Equation (2) to a deterministic ODE (probability flow) (e.g., Song et al., 2021). In this way, the only source of randomness is in the initial ensemble, which is then evolved using only the drift term in Equation (2). However, since deterministic sampling continuously reduces noise levels during reverse diffusion (Karras et al., 2022), the final samples may undersample $p_{0}$ and collapse to local modes. Moreover, another challenge with larger $D$ is that the fraction of initial ensemble members $\{\boldsymbol{x}_{i}\}$ within low-probability regions of $p_{0}$ quickly approaches 1, which results in inaccurate Monte Carlo estimates of the score function and resulting drift.
For certain classes of problems where some information about the covariance structure of the parameters is available a priori, we can employ a localization-type strategy. Specifically, the flexible form of Equation (1) allows us to choose the scaling matrix of the forward diffusion to be a localization matrix defined as the Cholesky-decomposed prior covariance. In other words, by defining a forward diffusion process that utilizes the prior covariance, we can “push” the diffusion process to stay “close” to the dominant subspace defined by the prior covariance. The implicit assumption is that the posterior will follow a similar covariance structure as defined by the prior. In essence, this localization strategy is a kind of linear dimensionality reduction strategy where we assume that we know the relevant subspace a priori (e.g., Webber & Morzfeld, 2023). As a result, this strategy becomes more useful as the dimensionality of the problem increases. In the examples presented here, we utilize an Ornstein-Uhlenbeck forward process for the forward diffusion process with drift and diffusion operators defined as
$\boldsymbol{b}_{t}$ $=\theta(\boldsymbol{\mu}-\boldsymbol{x}_{t}),$ (10a)
which evolves samples from $p_{0}$ to a multivariate Gaussian distribution $\mathcal{N}(\boldsymbol{\mu},\alpha\boldsymbol{\Sigma}_{\text{prior}})$ over a time scale $\theta>0$. Here, $\alpha$ is a scalar inflation factor for scaling the prior covariance in order to optionally inflate the initial ensemble. We chose the magnitude of $\theta$ such that as $t\rightarrow 1$ the process has converged to $\mathcal{N}(\boldsymbol{\mu},\alpha\boldsymbol{\Sigma}_{\text{prior}})$ in a distributional sense. The reverse diffusion process can then be determined by simply plugging $\boldsymbol{b}_{t}$ and $\boldsymbol{g}_{t}$ into Equation 2.
Figure 3: Marginal posterior samples and energy distance for different sampling methods for the 20-$d$ Bayesian linear regression. For the marginal plots (left and center), the outer ellipse represents the $3\sigma$ posterior uncertainty as computed by the analytic least squares solution. Energy distance (right) is computed relative to the least squares solution. Ensemble methods (ENS) without a localization strategy can successfully find the mode of the posterior, but significantly underestimate the posterior variance. When using an Ornstein-Uhlenbeck forward process (ENS+LOC) based on the prior distribution, posterior samples are drawn at a comparable accuracy to methods that require gradients of the log posterior. For ENS, we use a probability flow with ($\sigma_{\rm{min}}$, $\sigma_{\rm{max}})=(0.1,10)$ and $\boldsymbol{g}_{t}$ with the form in Table 1. The ENS+LOC parameters are specified in Table 2. BFGS samples for computing the energy distance were generated from the Hessian approximation.
We apply this strategy to a 20-dimensional Bayesian linear regression problem. The regression problem corresponds to estimating the coefficients, $\boldsymbol{x}\in\mathbb{R}^{20}$, of uniformly spaced B-splines placed in the columns of a linear operator $\boldsymbol{G}\in\mathbb{R}^{500\times 20}$, given a data vector $\boldsymbol{d}\in\mathbb{R}^{500}$ (Appendix C). The B-splines serve as smooth interpolating basis functions for reconstructing signals from noisy data. In this class of regression problems, it is common to enforce a prior constraint that the spline coefficients are correlated over a given length scale, which can be represented by a squared-exponential Gaussian Process kernel $k(x_{i},x_{j})=\exp\left[-d(x_{i},x_{j})^{2}/(2L^{2})\right]$, where $d(x_{i},x_{j})$ is the Euclidean distance and $L$ is the prescribed length scale (e.g., Jolivet et al., 2014; Webber & Morzfeld, 2023). Using this kernel to build the scale matrix $\boldsymbol{g}_{t}$ for the Ornstein-Uhlenbeck forward process (see Table 2 for full parameter specification), we are able to obtain samples from the target posterior distribution that have nearly the same accuracy as methods that utilize gradients of the log posterior, see Figure 3.
Table 2: Parameters for the 20-$d$ and 100-$d$ problems utilizing the Ornstein-Uhlenbeck forward process with localization based on the prior covariance. For both problems, we set $N_{\rm{ens}}=1000$.
Problem $\theta$ $\mu$ $\alpha$ $\Delta t_{\rm{init}}$ $N_{r}$
20-$d$ 0.1 0 $4^{2}$ 0.002 10
100-$d$ 0.05 0 $5^{2}$ 0.002 30
3.3 High-dimensional Problems
For this example, we aim to estimate a permeability field, $a(\boldsymbol{x})$, that leads to an observed pressure field, $P(\boldsymbol{x})$, in a porous medium resulting from a scalar field of fluid sources or sinks, $f(\boldsymbol{x})$. This system is governed by a partial differential equation (PDE) involving spatial gradients of $a(\boldsymbol{x})$ and $P(\boldsymbol{x})$ (see Appendix D). In this simplified experiment, we start with dense, noisy observations of $P(\boldsymbol{x})$ and assume that $f(\boldsymbol{x})$ is known. The inverse problem becomes estimating $a(\boldsymbol{x})$ such that the force balance described by the PDE in Appendix D is satisfied. This formulation allows for straightforward comparison with MCMC methods that leverage gradient information without having to backpropagate through a numerical solver. We model $a(\boldsymbol{x})$ as a linear combination of two-dimensional B-splines such that $a(\boldsymbol{x})=\boldsymbol{G}(\boldsymbol{x})\boldsymbol{m}$, where $\boldsymbol{G}(\boldsymbol{x})$ is the spline matrix and $\boldsymbol{m}\in\mathbb{R}^{100}$ (Appendix D).
Using the insights from the previous section, we apply our ensemble method with an Ornstein-Uhlenbeck forward process (Table 2) to sample from the posterior distribution of the spline coefficients, where the scaling matrix of the diffusion process is defined by the Cholesky-decomposed covariance matrix of the prior distribution. The prior covariance is constructed to encourage a smooth reconstruction of the permeability field. Compared with samples obtained from NUTS and MALA, our method successfully samples from the posterior distribution, although, similar to MALA, the samples exhibit slightly lower variance than those obtained with NUTS, see Figure 4.
Figure 4: Marginal posterior samples for select pairs of two-dimensional B-splines used to reconstruct a permeability field given observations of a pressure field. Our ensemble method with an Ornstein-Uhlenbeck localization process (ENS+LOC) is compared to NUTS and MALA. The ellipses in each plot correspond to the prior covariance (1-, 2-, and 3-$\sigma$). For this $100$-$d$ problem, the posterior samples obtained with ENS+LOC are similar to those obtained with the gradient-based NUTS and MALA.
This effect is likely driven by the relative weakness in the correlations (off-diagonal terms) of the prior covariance matrix used in the Ornstein-Uhlenbeck forward process, which in turn provides weaker constraints on the ensemble random walk during reverse diffusion. As a result, the likelihood ratio $p_{0}(\boldsymbol{x}^{\prime})/p_{\rm{is}}(\boldsymbol{x}^{\prime})$ in Equation (6) becomes dominated by a small subset of the ensemble members, causing the other members to drift towards that subset. We find this effect to be more pronounced when using scalar (isotropic) diffusion processes. Nevertheless, our ensemble method with the Ornstein-Uhlenbeck forward process is successfully able to sample a substantial region of the posterior distribution in high-dimensions without requiring gradients of the distribution.
4 Conclusions
In this study, we introduced the ensemble strategies that leverage the dynamics of particle ensembles to estimate a score function used to generate samples from a target distributions within a reverse diffusion process. This method, rooted in principles of importance sampling, is robust for low- to medium-dimensional problems where gradients of the posterior log probability or the forward model are unavailable. This is often the case for complex black box forward models where analytic gradients are intractable and automatic differentiation methods are not an option.
Our experiments, encompassing a hierarchy of problems, highlight the advantages of our method over traditional MCMC methods such as NUTS and MALA, even when these methods use gradient information. Particularly noteworthy is the ability to effectively navigate and sample from non-Gaussian distributions, which makes this approach an appealing choice especially in fields like geophysics and computational physics where such challenges are prevalent.
The ideas presented here are however not without challenges and warrant further investigation. The scalability to very high-dimensional problems remains a challenge that is rooted in the obvious curse of dimensionality. While our strategy of utilizing Ornstein-Uhlenbeck processes defined by prior information has shown promise, additional research is needed to optimize for such scenarios. In other Bayesian inference scenarios, priors may be relatively uninformative, which could limit the utility of the Ornstein-Uhlenbeck processes. Other works have approached this problem by training neural network models of the score, but those methods require taking gradients of the log posterior. It remains a challenge to build a completely gradient-free method that also works in high dimensions while remaining computationally efficient.
Software and Data
We implemented all algorithms and strategies presented in this work in JAX (Bradbury et al., 2018). In particular, we used the Equinox (Kidger & Garcia, 2021) and Diffrax (Kidger, 2022) packages to build a flexible implementation of forward-reverse diffusion systems.
Broader Impact
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Anisotropic Forward Diffusion
For higher-dimensional problems where some information on the covariance structure of the target distribution is available (which is common for geophysical systems, due to physical constraints), Ornstein-Uhlenbeck processes with non-zero drift coefficients and with diffusion informed by the known covariance structure (Equation 3.2) can significantly improve the accuracy of the reverse diffusion process. For scalar Brownian motion, the samples will take random walks in each dimension with equal variance. For higher-dimensional problems, this behavior can cause many samples to move towards low probability regions of the target distribution. With the Ornstein-Uhlenbeck process, the random walks are constrained to the covariance structure imposed by $\boldsymbol{g}_{t}$, which for certain problems can keep samples in the relevant regions of the probability space to improve the estimate of the Föllmer drift (Figure 5). In practice, we also tune $\theta$ to control the relative importance of $\mu$ during the reverse diffusion. For $\theta=1$, the samples will stay close to $\mu$, which can be un-desirable when the target mean is expected to differ non-negligibly from $\mu$. For the higher-dimensional experiments in this work, we find $\theta=0.1-0.2$ to provide the best accuracy.
Appendix B Energy distance
The energy distance $\epsilon$ is related to the covariance distance between two random vectors $\boldsymbol{X}$ and $\boldsymbol{Y}$ and is defined as (Székely & Rizzo, 2013):
where $\|\cdot\|_{p}$ is the $p$-norm and $\boldsymbol{X}^{\prime}$ and $\boldsymbol{Y}^{\prime}$ are random permutations of $\boldsymbol{X}$ and $\boldsymbol{Y}$, respectively. As such, the estimate for $\varepsilon(\boldsymbol{X},\boldsymbol{Y})$ can be repeated for multiple realizations of $\boldsymbol{X}^{\prime}$ and $\boldsymbol{Y}^{\prime}$ such that the mean value represents a more robust estimate. Here, we use the Python package dcor package to compute $\varepsilon(\boldsymbol{X},\boldsymbol{Y})$ (Ramos-Carreño & Torrecilla, 2023). As this estimate can vary depending on the size of the vectors $\boldsymbol{X}$ and $\boldsymbol{Y}$, comparing the samples generated by our method for different ensemble sizes requires multiple runs of our method with different initial random ensembles until a vector of a fixed size $N_{\text{total}}$ is filled, where $N_{\text{total}}$ is several multiples of $N_{ens}$. Overall, we find the energy distance to be a robust way of comparing samples from arbitrary probability distributions, and we find that using a $p$-norm with $p=1$ or 1.5 can mitigate the effects of outliers in $\boldsymbol{X}$ and $\boldsymbol{Y}$.
Appendix C Bayesian linear regression
For the linear model used in Section 3.2, the matrix $\boldsymbol{G}$ is constructed from 20 equally-sized B-splines uniformly distributed along some spatial dimension (Figure 6a.) (Hetland et al., 2012).
@@ Appendix C Bayesian linear regression
The spline coefficients, represented by the vector $\boldsymbol{x}$, are assumed to be spatially correlated following the prior covariance structure shown in Figure 6b, which is constructed from the squared-exponential Gaussian Process kernel as described in Section 3.2. In this experiment, the B-splines are defined within a spatial domain of $[-1,1]$, and the prior length scale $L$ is set to $0.5$.
The posterior distribution for the spline coefficients follows from Bayes’ Theorem (ignoring the normalization constant):
where the first distribution on the right-hand side is the likelihood distribution and the second distribution is the prior distribution. For a given data vector $\boldsymbol{d}=\boldsymbol{Gx}+\boldsymbol{\epsilon}$, where $\boldsymbol{\epsilon}\sim\mathcal{N}(\boldsymbol{0},\sigma_{d}^{2}\boldsymbol{I})$ is normally distributed noise with variance $\sigma_{d}$, and a multivariate Gaussian prior, $p(\boldsymbol{x})\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{\Sigma}_{\rm{prior}})$, we can derive an analytic solution for $p(\boldsymbol{x}|\boldsymbol{d})$ as the least-squares estimate (Tarantola, 2005):
@@
Together, $\hat{\boldsymbol{x}}$ and $\boldsymbol{\Sigma}_{\hat{\boldsymbol{x}}}$ are used to construct a reference multivariate Gaussian distribution to compare the various sampling methods. For this experiment, we generate $\boldsymbol{d}$ by sampling a reference $\boldsymbol{x}$ from a unit Gaussian distribution and multiplying by $\boldsymbol{G}$. We additionally add noise with a length scale of 0.05 and a variance of $\sigma_{d}=2.0$ to obtain the final $\boldsymbol{d}$ used in the experiments.
Appendix D Darcy Flow
For a spatial domain defined by coordinates $\boldsymbol{x}$, a permeability field, $a(\boldsymbol{x})$, and a scalar field of fluid sources or sinks, $f(\boldsymbol{x})$, the following PDE can be derived for a pressure field $P(\boldsymbol{x})$ following the derivation of Darcy flow (e.g., Garbuno-Inigo et al., 2020):
@@ Appendix D Darcy Flow
where $\mathcal{D}$ represents the interior of the modeling domain and $\partial\mathcal{D}$ represents the boundary. We prescribe $\mathcal{D}$ to span the domain $[0,1]^{2}$ and assume that all fields have normalized units. We solve the above equations for a unit square mesh with Dirichlet boundary conditions using firedrake (Ham et al., 2023). The ground truth $a(\boldsymbol{x})$ is modeled as $a(\boldsymbol{x})=a_{0}e^{\theta(\boldsymbol{x})}$, where $a_{0}=0.6$ and and $\theta(\boldsymbol{x})$ is a random Gaussian field with a length scale of 0.15. The simulated fields are shown in Figure 7.
@@
In other words, we aim to estimate $a(\boldsymbol{x})$ that minimizes the force balance residual, which is a strategy widely used in physics-informed machine learning (e.g., Raissi et al., 2019). This approach removes the need for backpropagating through a numerical solver and facilitates comparison of our approach with methods that require gradients through the log posterior of $\theta(\boldsymbol{x})$. Each spatial field is resampled to a $100\times 100$ square grid such that the gradients in Equation (15a) can be evaluated with finite differences.
However, the uniform grid still results in 10,000 grid cells. In order to reduce the dimensionality of the spatial fields for the inverse problem, we represent the fields as a linear combination of $100$ 2D B-splines distributed uniformly in both the horizontal and vertical directions (Figure 8). The two-dimensional B-splines are simple outer products of the one-dimensional B-splines used in the linear regression problem and can decompose spatial fields in a manner similar to wavelet transforms. Splines of multiple length scales can be combined in order to accurately represent multiscale fields. Here, we simply use a single length scale for our analysis.
@@
Each spline is evaluated on the same $100\times 100$ square grid as the pressure and source fields, flattened to a column vector, and placed into a column of a matrix $\boldsymbol{G}\in\mathbb{R}^{10000\times 100}$. The posterior distribution for the spline coefficient vector, $\boldsymbol{m}$, is (here we use $\boldsymbol{m}$ to avoid confusion with the coordinate space defined by $\boldsymbol{x}$):

### v2

@@ 4.3 Moderate-dimensional Problems
As we expect the intrinsic variance of the Monte Carlo estimation of the score function to increase rapidly with increasing dimensionality, it is useful to investigate a moderate-dimensional problem ($D<10$) to characterize the sampling behavior of our proposed methods and the utility of our variance reduction strategies. To that end, we formulate a 20-dimensional Bayesian linear regression problem corresponding to estimation of the coefficients, $\boldsymbol{x}\in\mathbb{R}^{20}$, of uniformly spaced B-splines placed in the columns of a linear operator $\boldsymbol{G}\in\mathbb{R}^{500\times 20}$, given a data vector $\boldsymbol{d}\in\mathbb{R}^{500}$. The B-splines serve as smooth interpolating basis functions for reconstructing signals from noisy data (Figure 7a.) [34].
@@
In this class of Bayesian regression problems, it is common to enforce a prior constraint that the spline coefficients, represented by the vector $\boldsymbol{x}$, are correlated over a given length scale, which can be represented by a squared-exponential Gaussian Process kernel $k(x_{i},x_{j})=\exp\left[-d(x_{i},x_{j})^{2}/(2L^{2})\right]$, where $d(x_{i},x_{j})$ is the Euclidean distance and $L$ is the prescribed length scale (e.g., [29, 35]). In this experiment, the B-splines are defined within a spatial domain of $[-1,1]$, and the length scale $L$ is set to $0.5$.
With this setup, the aim is obtain samples from the posterior distribution of the spline coefficients given a set of (potentially noisy) observations. For a given data vector $\boldsymbol{d}=\boldsymbol{Gx}+\boldsymbol{\epsilon}$, where $\boldsymbol{\epsilon}\sim\mathcal{N}(\boldsymbol{0},\sigma_{d}^{2}\boldsymbol{I})$ is normally distributed noise with variance $\sigma_{d}$, and a multivariate Gaussian prior, $p(\boldsymbol{x})\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{\Sigma}_{\rm{prior}})$, we can derive an analytic solution for $p(\boldsymbol{x}|\boldsymbol{d})$ as the least-squares estimate [36]:
@@
Together, $\hat{\boldsymbol{x}}$ and $\boldsymbol{\Sigma}_{\hat{\boldsymbol{x}}}$ are used to construct a reference multivariate Gaussian distribution which we can use to assess the accuracy of various sampling methods. For this experiment, we generate $\boldsymbol{d}$ by sampling a reference $\boldsymbol{x}$ from a unit Gaussian distribution and multiplying by $\boldsymbol{G}$. We additionally add noise with a length scale of 0.05 and a variance of $\sigma_{d}=2.0$ to obtain the final $\boldsymbol{d}$ used in the experiments.
4.3.1 Sampling the Posterior Distribution
For low-dimensional problems, the linear operator $\boldsymbol{b}_{t}$ and the scaling matrix $\boldsymbol{g}_{t}$ in Equation (1) can be replaced with scalar terms such that the forward diffusion process evolves with equal variance in each dimension. As mentioned in Section 3.2.4, as the dimensionality increases, isotropic random walks can cause ensemble members to move to areas of low probability, similarly to the standard Metropolis-Hastings algorithm. One commonly used strategy in score-based sampling is to convert the SDE in Equation (2) to a deterministic ODE (probability flow) [5]. In this way, the only source of randomness is in the initial ensemble, which is then evolved using only the drift term in Equation (2). However, since deterministic sampling continuously reduces noise levels during reverse diffusion [31], the final samples may undersample $p_{0}$ and collapse to local modes. Moreover, another challenge with larger dimensionality is that the fraction of initial ensemble members $\{\boldsymbol{x}_{i}\}$ within low-probability regions of $p_{0}$ quickly approaches 1, which results in inaccurate Monte Carlo estimates of the score function and resulting drift.
As discussed in Section 3.2.4, we can utilize an Ornstein-Uhlenbeck forward process to constrain the trajectories of the samples during reverse diffusion. Here, we construct the scale matrix $\boldsymbol{g}_{t}$ directly from the prior covariance (see Table 3 for full specification of the Ornstein-Uhlenbeck parameters). This variance reduction strategy allows us to obtain samples from the target posterior distribution that have nearly the same accuracy as methods that utilize gradients of the log posterior (as measured by the energy distance between the generated samples and the analytic posterior distribution; see Figure 8). While our method with a simple isotropic Gaussian diffusion term can successfully locate the mode of the posterior distribution, the sample variance is too low, leading to a high energy distance. MALA is able to increase the variance of the samples, albeit lower than that expected from the true posterior distribution. We note that MALA with fine-tuned Langevin dynamics could provide marginal improvement but would still be limited by its use of isotropic Brownian motion, which places it under the same limitations as our method with an isotropic diffusion term. Our method with an Ornstein-Uhlenbeck forward process performs nearly as well as NUTS, which requires gradients of the posterior distribution to fully explore the probability space. Moreover, compared to other gradient-free algorithms, our method achieves higher accuracy with significantly fewer forward evaluations, $N_{\text{fev}}$.
Figure 8: Comparison of sampling methods for 20-D Bayesian linear regression. Left: Marginal posterior samples versus $3\sigma$ analytic uncertainty. Right: Energy distance vs. number of forward evaluations (blue-green=gradient-free, orange=gradient-based). Basic Ensemble method (ENS) underestimates variance, while other gradient-free methods (Particle Monte Carlo - PMC, Ensemble Slice Sampling - ESS) are computationally costly. ENS+LOC, using a prior-informed Ornstein-Uhlenbeck process, achieves accuracy comparable to Ensemble Kalman Sampling (EKS) and gradient-based methods while requiring fewer forward evaluations without gradients.
Table 3: Parameters for the 20-$d$ and 100-$d$ problems utilizing the Ornstein-Uhlenbeck forward process with localization based on the prior covariance.
Problem $\theta$ $\mu$ $\alpha$ $\Delta t_{\rm{init}}$ $N_{\rm{ens}}$ $N_{r}$
20-$d$ 0.1 0 $4^{2}$ 0.002 1000 10
100-$d$ 0.05 0 $5^{2}$ 0.002 2000 30
4.4 High-dimensional Problems
For this example, we aim to estimate a permeability field, $a(\boldsymbol{x})$, that leads to an observed pressure field, $P(\boldsymbol{x})$, in a porous medium resulting from a scalar field of fluid sources or sinks, $f(\boldsymbol{x})$. This system is governed by the following PDE for the pressure field $P(\boldsymbol{x})$ following the derivation of Darcy flow (e.g., [17]):
@@ 4.4 High-dimensional Problems
where $\mathcal{D}$ represents the interior of the modeling domain and $\partial\mathcal{D}$ represents the boundary. We prescribe $\mathcal{D}$ to span the domain $[0,1]^{2}$ and assume that all fields have normalized units. We solve the above equations for a unit square mesh with Dirichlet boundary conditions using firedrake [37]. The ground truth $a(\boldsymbol{x})$ is modeled as $a(\boldsymbol{x})=a_{0}e^{\theta(\boldsymbol{x})}$, where $a_{0}=0.6$ and $\theta(\boldsymbol{x})$ is a random Gaussian field with a length scale of 0.15. The simulated fields are shown in Figure 9.
@@
In other words, we aim to estimate the permeability field that minimizes the force balance residual for an observed pressure field, which is a strategy widely used in physics-informed machine learning [38]. This approach removes the need for backpropagating through a numerical solver and facilitates comparison of our approach with methods that require gradients through the log posterior of $\theta(\boldsymbol{x})$. Each spatial field is resampled to a $100\times 100$ square grid such that the gradients in Equation (17a) can be evaluated with finite differences.
4.4.1 Spline Representation
We represent the $\theta(\boldsymbol{x})$ field as a linear combination of $100$ 2D B-splines distributed uniformly in both the horizontal and vertical directions (Figure 10). The two-dimensional B-splines are simple outer products of the one-dimensional B-splines used in the linear regression problem in Section 4.3 and can decompose spatial fields in a manner similar to wavelet transforms. Splines of multiple length scales can be combined in order to accurately represent multiscale fields. Here, we simply use a single length scale for our analysis.
@@ 4.4.1 Spline Representation
Each spline is evaluated on the same $100\times 100$ square grid as the pressure and source fields, flattened to a column vector, and placed into a column of a matrix $\boldsymbol{G}\in\mathbb{R}^{10000\times 100}$. Overall, while the spline representation significantly reduces the dimensionality of the inverse problem, it still provides a bridge between the low- and moderate-dimensional problems investigated in the previous sections and higher-dimensional problems common in physical applications. Moreover, the spline representation admits a physically-motivated prior covariance matrix, which we now discuss.
The posterior distribution for the spline coefficient vector, $\boldsymbol{m}$, is (here we use $\boldsymbol{m}$ to avoid confusion with the coordinate space defined by $\boldsymbol{x}$):